OpenStudy (anonymous):

Solve (9x + 2) – 3(x – 4) = 6.

OpenStudy (anonymous):

x = 4/3 x = –4/3 It is an identity and is true for all real numbers. There is no solution. answers

OpenStudy (anonymous):

6x+2+12=6

OpenStudy (anonymous):

6x=-8

OpenStudy (anonymous):

x=-4/3

OpenStudy (anonymous):